OpenStudy (ell):

Medal to helper!! If x^n=1 for all real numbers x(= crossed equals sign)0, then what does n equal?

jimthompson5910 (jim_thompson5910):

By definition, x^0 = 1. The variable x can be any real number but 0. In other words, we can't say 0^0 = 1, but we can say something like 2^0 = 1 or 3^0 = 1, etc.

OpenStudy (ell):

Well the choices are: a. -1 b. 1 c. 0 d. x Would it be B or can it go into negatives?

jimthompson5910 (jim_thompson5910):

compare x^n = 1 with what I wrote: x^0 = 1

jimthompson5910 (jim_thompson5910):